Go to Planning Wiki - MB Wiki - MB Web - MB Members - MB Meetings - MB Minutes - MB Action List

MB Older Actions

  • A.Aimar will ask information about the CERN reliability data during the power cut.

When CERN was down, no new SAM tests were run. This can be seen for CERN in gridview[1], where there is a large gap that day from 03:22:55 to 30/05/2008 14:49:53.

The last test CERN had (at 3:22) was GREEN, so during the whole period, it was considered GREEN by Gridview – thus availability was 100%

After the fact, two downtimes were entered for CERN : 06:36-08:00 and another from 08:13-10:00. Gridview only takes into consideration downtimes entered 2 hours BEFORE an intervention. Therefore, reliability was 100%.

  • H.Marten and Y.Schutz agreed to verify and re-derive the correct data and the normalization factors applied (at FZK for instance). And compare the local accounting, APEL WLCG accounting and the ALICE MonALIsa accounting.

Done.

  • 30 Apr 2008 - Sites send to H.Renshall plans for the 2008 installations and what will be installed for May and when the full capacity for 2008 will be in place.

The issue was discussed at the LHCC mini review and J.Gordon received the updated information. J.Gordon will pass the information to H.Renshall.

  • 1 June 2008 - J.Templon will summarize the proposal of the interim solution for treating Operations Alarms until a GGUS-based solution is available.

Done.

  • 31 May 2008 - OSG Tier-2 should report availability and reliability data into SAM and GOCDB; so that the monthly Tier-2 report can include their information.

To remove, New milestones have been already added to the HL Milestones dashboard.

Done. The agenda above is considered accepted.

  • 16 May 2008 - I.Bird should ask information to CERN about the plans of the Accelerator for 2009.

  • 24 May 2008 - A.Aimar will set up a page for the site operations alerts.
Update: The page for Operations Alarms is here: https://twiki.cern.ch/twiki/bin/view/LCG/OperationsAlarmsPage

  • 18 May 2008 - Sites should send the emails address where they are going to receive operations alarms. Sites should confirm that they can achieve this.

Obsolete. Replaced by the “Operations Alarms Page” action.

A new page for alert emails for the sites, not the same as the contact pages. This is a separate page to be prepared. Is not the same and the contact page already prepared (https://twiki.cern.ch/twiki/bin/view/LCG/TierOneContactDetails). Should be a different list only for alerts and only for a few people from the Experiments. OperationsAlarms

  • 31 March 2008 - ALICE, ATLAS and CMS should provide the read and write rates that they expect to reach. In terms of clear values (MB/sec, files/sec, etc) including all phases of processing and re-processing.

Not Done. Next week we will close this item and address specific issues in the future. It was agreed that the Experiments provide the same information in the format used by LHCb

Here is the example from LHCb: https://twiki.cern.ch/twiki/pub/LCG/GSSDLHCB/Dataflows.pdf

  • 31 March 2008 - OSG should prepare site monitoring tests equivalent to those included in the SAM testing suite.
    J.Templon and D.Collados will verify this equivalence and report to the MB as it was done for NDGF.

Done in the previous MB meeting. The OSG tests are described here: http://rsv.grid.iu.edu/documentation/help/.

The proposed new list of critical tests is available here: https://twiki.cern.ch/twiki/bin/view/LCG/OSGCriticalProbes#Proposed_Critical_Probes_for_OSG

  • 6 May 2008 - Sites should confirm to the MB that they will define an email digitally signed (?) for their alarm system submission.

Replaced by the action to fill in the wiki page.

  • 16 May 2008 - Each Experiment proposes 4 users who can raise alarms at the sites and are allowed to mail to the sites alarm mailing list.

Replaced by the action to fill in the wiki page.

  • 9 May 2008 - Sites should send comments about the New HL Milestones.
Done.

  • 18 Mar 2008 - Sites should propose new tape efficiency metrics that they can implement, in case they cannot provide the metrics proposed.

Done. The sites that cannot implement the metrics have reported to the reasons and a HL milestone is introduced to track the situation. Here is the link to the wiki page: https://cern.ch/twiki/bin/view/LCG/MssEfficiency

FZK: H.Marten explained that those metrics at FZK still need to be implemented and will not be ready for CCRC08-May. They could publish other metrics in the wiki page but he would first make sure that the data is correct.
ASGC: Have received the scripts from CERN in order to extract the metrics from their CASTOR installation. They still have to modify them in order to make them work with their specific setup.
IN2P3 was not present but F.Hernandez sent an email after the meeting:FR-CCIN2P3 is one of the sites not yet publishing this information. For your information, we are currently collecting the data for extracting the proposed metrics as closely as possible. The work on the tools to automate the extraction of the relevant information and generation of the metrics was more time-consuming than what we initially anticipated. We expect to have this initial phase finished by the end of this week and start publishing the available data in the wiki by the end of next week.

  • 18 April 2008 - I.Bird and A.Aimar will propose new milestones to the Management Board.

First proposal on the 29.4.2008. WIll be discussed at the F2F MB.

  • 29 Feb 2008 - A.Aimar will verify with the GridView team the possibility to recalculate the values for BNL.

Done. The BNL values are recalculates and will be included in the the report of the Tier-1 reliability report.

  • 15 Apr 2008 - A.Aimar will distribute the VO-specific milestones for March 2008.

Done.

  • 15 Apr 2008 - A.Aimar- A note should be added to the Tier-2 Reliability reports indicating that the US sites are not yet reporting availability and reliability data.

Done. The US Tier-2 sites have not available “n/a” instead of the 0% as before.

  • 18 Apr 2008 - M.Ernst should clarify the situation with the ATLAS sites that are not providing and SRM interface. And how availability reliability is reported.

Done. M.Ernst explained that there is the commitment of all US ATLAS sites to have an SRM interface by and of May 2008.

  • 29 Feb 2008 - A.Aimar will verify why the reliability values for the Tier-2 sites seems incorrect (being lower than availability).

Done. Reliability is lower than availability because availability was considering the “unknown” status as “available” in order not to penalize the sites when SAM has problems. The algorithm is going to be modified to fix this inconsistency. For the availability calculations if “unknown” is between “downtime” periods it will be considered “unknown = downtime” and not “unknown = available”.

Done.

  • 26 Feb 2008 - Discuss/agree on a milestone about providing tape efficiency metrics should be introduced for the Tier1-s sites, before CCRC08 in May.

Done. The MB Members agreed that a milestone should be added in order to have metrics about tape efficiency at all Tier-1 sites. In order to be ready for the CCRC08-May the metrics should be in place by end of April 2008. The metrics to produce are the ones in the Tape Efficiency Wiki or equivalent. If sites can provide more metrics is even better.

  • 26 Feb 2008 - A.Aimar verifies the values of the GridView reliability calculations for NL-T1.
Done.

  • 22 Feb 2008 - Experiments should provide feedback on the priorities of the current storage issues. See this link.

Done. Input from ATLAS, CMS and LHCb received.

To be replaced by a milestones for the sites to prepare tape efficiency metrics.

  • 13 Feb 2008 - A.Aimar will ask for the Sites Reliability Reports January 2008. Sites should complete them by end of the week.
Done.

  • 24 Jan 2008 - Sites should update the High Level Milestones. A.Aimar will remind it via email during the week.
Done.

  • 5. Feb.2008 - I.Bird will find a speaker for the Experiments Status at the LHCC Referees Meeting.
Removed. The Experiments will have each a 10' presentation.

  • 21 Jan 2008 - The LCG Office should define where (a web area, wiki, share point?) the Sites can upload their statistics about their tape storage performance and efficiency.

Done. Individual pages for publishing tape efficiency data, for each Tier-0 and Tier-1 Site, are now available.

  • 18 Dec 2007 - LHCb should nominate who is responsible for the benchmarking of their applications on the machines made available by the HEPiX Benchmarking Working Group.

Done. Ph.Charpentier will be included in the mailing list until LHCb nominates someone.

  • 11 Jan 2008 - H.Meinhard will distribute the information about HEP Benchmarking to the contacts for benchmarking in the Experiments.

Done

  • 10 Jan 2008 - A.Aimar will ask for the Site Reports for December 2007. Experiments should review their specific SAM tests and see whether the GridView summary is correct.

Done

  • 14 Jan 2008 - A.Aimar will communicate these names of the OPN Reps. to D.Foster and ask him which are the Tier-1 sites that are not properly represented in order to address them directly.

Done

  • 9 Jan 2008 - T.Bell will distribute to the MB a set of metrics that could be also measured by all Tier-1 Sites in order to describe in a uniform way the performance of the tape storage systems.

Done

  • 8 Jan 2008 - Ph.Charpentier agreed to distribute the 2008 LHCb needs for each site to the MB mailing list and to the LHCb national representatives.

Done at the CCRC Meeting on the following Thursday.

  • 8 Jan 2008 - ALICE and ATLAS will present at the MB F2F in January how they intend to solve the issues caused by tiny files.

Done.

  • 8 Jan 2008 - K.Bos had prepared a document describing the ATLAS requirements for the CCRC. J.Shiers will distribute it to the CCRC list.

Done already before Christmas. And will be presented to the CCRC on the following Thursday.

  • 18 Dec 2007 - Experiments should nominate who is responsible for the benchmarking of their applications on the machines made available by the HEPiX Benchmarking Working Group.

Done for ALICE, ATLAS CMS. Remains for LHCb.

ALICE: Peter Hristov

ATLAS: Alessandro De Salvo and Franco Brasolin for technical help.

CMS: Gabriele Benelli

  • 11 Dec 2007 - L.Dell’Agnello, F.Hernandez and G.Merino prepare a questionnaire or a check list for the Experiments in order to collect the Experiments requirements in a form suitable for the Sites.

Done. The follow-up from the Experiments is for the CCRC08 F2F Meeting on Thursday.

  • 30 Nov 2007 - The Tier-1 sites should send to A.Aimar the name of the person responsible for the operations of the OPN at their site.

Considered completed with the information received so far.

Received information from TW-ASGC (Min Tsai, Aries Hung), FR-CCIN2P3 (Jerome Bernier), IT-INFN (Stefano Zani), RAL (Robin Tasker), DE-KIT (Bruno Hoeft), PIC (G. Merino)

* 21 Oct 2007 - Sites should send to H.Renshall their resources acquisition plans for CPU, disks and tapes until April 2008

Done. NDGF and NL-T1 should send to H.Renshall and S.Foffano an estimate about the delivery of their 2008 capacity.

Update: NL-T1 send a mail with November 2008 as estimate for making available the 2008 capacity.

  • L.Robertson will prepare a note reporting and explaining the current definitions of “availability” and “reliability” wrt. the definitions in the MoU.
Done.

  • 23 October 2007 - D.Barberis agreed to clarify with the Reviewers the kind of presentations and demos that they are expecting from the Experiments at the Comprehensive Review.
Done.

  • 30 October 2007 - I.Fisk will ask R.Pordes for information about glexec with other OSG VOs and sites. And whether OSG has already reviewed the code of glexec from the security point of view.
Done.

  • 6 November 2007 - J.Gordon and A.Aimar will send a proposal to the MB for an agreement on the acceptance of pilot jobs and glexec.
Done by L.Robertson

  • 18 Sept 2007 - Next week D.Liko will report a short update about the start of the tests in the JP working group.

  • 21 Sept 2007 - D.Liko sends to the MB mailing list an updated version of the JP document, including the latest feedback.

I.Bird agreed to report to the Management Board about the progress of the Job Priority working group.

  • 16 October 2007 - Sites should send the pointers to their documents about 24x7 and VO Boxes to A.Aimar. A.Aimar will prepare a protected web area for confidential documents of the LCG Management Board.

Done. The MB LCG Private Web Area (link) is available and the access is limited to the members of the MB mailing list. Sites can upload their documents. In case of problems contact A.Aimar.

  • 10-July 2007 - INFN will send the MB a summary of their findings about HEP applications benchmarking.

Done.

  • 1 Oct 2007 - A.Aimar will distribute to the MB the Sites Reliability table for September 2007 and Sites will respond at the F2F Meeting in October.

Done.

  • 25 September – MB Members send feedback on the new GridView computation algorithm.

Done.

  • 25 Sept 2007 - J.Gordon distributes the Grid Security and Grid Operation policy documents to approve at the F2F MB meeting
Done.

  • 31 Dec 2007 - L.Robertson will present the agreement(s) to propose to the different supplier of external software for the LCG.
Postponed to end of 2007.

  • 25 Sept 2007 - I.Bird will propose speakers for the Tier-1 and Tier-2 presentations at the LHCC Comprehensive Review.

Removed. The Agenda for the LHCC Comprehensive Review is anyway going to be discussed further in this meeting.

  • 25 Sept 2007 - A.Aimar finds out the information about the VO-specific tests and adds it to the SAM wiki page (contacts, test documentation, etc).

  • 28 August 2007 - SAM tests in July 2007: ALICE on all tests, ATLAS@IN2P3 and CMS@INFN will investigate and report on the VO-specific results.

ALICE not done.

CMS@INFN Done. M.Kasemann explained that CMS now reviewed the critical tests because there were some tests failing at CNAF. A.Sciabá should have fixed that.

ATLAS@IN2P3 Done. By email F.Hernandez explained that ATLAS and IN2P3 have investigated the issue and it will be solved. The failures were due to the attempt of some tests to create some files in an area where they did not have write permissions.

  • 18 Sept 2007 – Sites and Experiments will send to J.Shiers the name of their representative in the CCRC Coordination team.

Done.

  • 15 Sept 2007 - The MB members should send feedback and comments to A.Aimar about the changes to the QR report.
Done.

  • 12 July 2007 - Progress of the Job Priorities working group and the progress of the working solution for ATLAS. D.Liko has been asked to provide a summary, on status and plans of the JP Working Group, which will be distributed to the MB by I.Bird.
Done.

  • 31 August 2007 - J. Shiers to organise discussions at the WLCG Workshop in Victoria on compatibility and SRM implementation issues and the US planning process.
Done.

  • 31 August 2007 - L. Robertson to make an SRM Management meeting written proposal following discussion with J. Shiers and F. Donno.
Removed. Proposal distributed by L.Robertson but not accepted. See Section 1.2 above.

  • 12 July 2007 - L.Robertson will appoint assessors to review the equivalence of the OSG tests and the WLCG test set agreed in April 2006.
Done.

  • 10 August 2007 - Sites and Experiments that disagree with the status of the HL milestones in the table below should send their feedback to A.Aimar before the 10 August 2007
Done.

  • 31 July 2007 – J.Shiers – Analysis of site scheduled and unscheduled interruptions, showing relative impact at sites, frequency and duration of different causes, etc.
Obsolete. J.Shiers presented the analysis of the interruptions in a previous meeting and this will become a standing activity that will be discussed at the Workshop.

  • 24 July 2007 - CNAF and LHCb to specify additional milestones for StoRM to add to the SRM Roll-out plan (i.e. extending milestones SRM-7 and SRM-17 to include testing of StoRM).
Done. Milestone SRM-7A added to the SRM2 Roll-out Plan.

  • 24 Jul 2007 - L.Robertson will ask via email to the Experiments to provide their update to the requirements until 2012.
Done. The status of the requirements is discussed in this meeting.

  • 6 July 2007 - Tier-0 and Tier-1 sites complete the Site Availability Reports for June 2007.

Done.

  • 10 July 2007 - L.Robertson asked that, for next week, F.Donno presents a detailed SRM roll-out plan with all the activities that have to be executed and deadlines for sites, developers and experiments.
Done. The SRM roll-out plan will be presented and discussed at this MB meeting.

  • 3 July 2007 - I.Bird and C.Grandi will report on the structure and leadership reformed Job Priorities Working Group.

Done.

I.Bird reported that the Working Group has been reformed. For the moment the group is discussing a short-term solution for ATLAS, not the long-term more general solutions. It is led by D.Liko and already met on a couple of occasions.

C.Grandi added that all focus is on the short-term solution for ATLAS. There was a presentation on the “gpbox” solution, but this will not be discussed until the working solution for ATLAS is implemented and available on the PPS.

  • 29 May 2007 - A.Aimar and L.Robertson circulate follow-up milestones on the VOs Top 5 Issues to the MB.
Done.

  • 15 May 2007 - J.Gordon agreed to make a proposal on how to also store non-grid usage in the APEL repository.
Done.

  • 22 May 2007 - I.Fisk will circulate to the MB the description of the CMS SAM tests.
Done.

  • 22 May 2007 - L.Robertson will distribute a proposal that had been prepared after the MB presentation on benchmarking in March.
Done. Proposal distributed and a talk is scheduled for the F2F MB meeting next week.

  • 15 Mar 2007 - CMS and LHCb should send to C.Eck their requirements until 2011.
Action removed. C.Eck proposed to have this item removed because, by July 2007, a new update of the requirements (until 2012) is already scheduled.

  • 24 Apr 2007 - In order to be in sync of the meetings of the Overview Board L.Robertson and A.Aimar will propose new dates for the preparation of future QR Reports.
Done. The MB agreed to move forward of one month the QR reports for 2007, to July, October and January.

  • 3 Apr 2007 - A.Aimar will update the targets for all four LHC experiments and distribute it to the MB.

  • 10 Mar 2007 - ALICE and LHCb should send to A.Aimar their targets for 2007, in a similar format to those specified for ATLAS and CMS.
Done. Alice and LHCb sent their targets for 2007.

  • 27 Feb 2007 - H.Renshall agreed to summarize the experiments; work for 2007 in an “Overview Table of the Experiments Activities.
Done. The calendars by experiment and site are already available on the SC4 Experiments Plans Wiki.

  • 16 Mar 2007 - Tier-0 and Tier-1 sites should send to the MB List their Site Reliability Reports for February 2007
Done. Summary presented at this MB meeting.

  • 31 Mar 2007 - Experiments to provide very long term (2010-2020) estimates of their computing needs at CERN
Cancelled. L.Robertson had discussed the issue with J.Engelen. Beyond 2009 there are several options about the evolution of the LHC, but nothing has been decided yet and so computing estimates must remain speculative.

  • 22 Feb 2007 - Experiments (ALICE and LHCb in particular) should verify and update the Megatable (the Tier-0-Tier-1 peak value in particular) and inform C.Eck of any change.
Done. C.Eck reported that LHCb’s values are now correct. The values from CMS and ALICE still need to be updated.

  • 6 Mar 2007 - L.Robertson will organize a meeting with someone from the accelerator to discuss the planning until 2020.
Done. L.Robertson discussed the issue with J.Engelen. Beyond 2009 there are several possible open options about the evolution of the LHC.

J.Engelen will speak to the DG about the evolution of the CERN computing facilities. The issue could be postponed until it is clear how the LHC will be upgraded in the future. In the mean time one could base the requirements on evolutionary estimations, based on past experience.

  • 27 Feb 2007 - MB Members should review and send feedback on the note distributes by L.Robertson on the process for reporting site capacity and usage data.
Done. Feedback received and an extended version is discussed in this meeting.

  • 9 Feb 2007 - All Tier-1 Sites should send to the MB their Site Reliability Report for December 06 and January 07.
Done. A short summary is presented later in this meeting.

  • 20 Feb 2007 - MB Members should send feedback to L.Robertson on the list of products and on the kind of agreement to reach with the external software providers.
Done. L.Robertson will send a proposal on how to proceed with the different suppliers of software.

  • 31 Jan 2007 - All Tier-1 Sites should send to H.Renshall their procurement plans for 2007.
Done. H.Renshall said that he has the data he needs and that should to be verified by the sites.

  • 5 Feb 2007 - All Tier-1 Sites should send to the MB their plans for the implementation of job priorities.
Done. An update is done with the roundtable in MB meeting 06/02/2007.

  • 5 Feb 2007 - L.Robertson will send a starter list of software involved to the MB.
Done.

  • 19 December - J.Shiers and H.Renshall will report on the progress on the definition of targets and milestones for 2007 at the LCG ECM meeting.

Action closed. Initial input from experiments received. A first proposal of the 2007 targets should be available early January 2007.

  • 25 Jan 2007 - F.Donno should report to the MB (by email) the list of people that will participate to the GSSD working group.

Done. The list has been distributed to the MB list and a few comments have circulated.

* 22 Jan 2007 - J.Templon should distribute pointers to all information and documentation that can help sites to implement job priorities on their batch systems.

Done. R.Starink reported that the information is available in the presentation given at the WLCG Workshop. J.Templon distributed an email to the MB mailing list with further information.

  • 22 January – All Tier-1 sites should send their site report on December 2006 reliability.
Replaced by a new action including also January 2007.

  • 15 Jan 2007 - D.Foster will form a group discussing the network setup and performance needed according to the Megatable values (for T1-T1 links on the OPN, etc).

Done.

  • 19 Dec 2006 - The proposal to ALICE is to consider, as in the TDR, a value of 10**6 for the ALICE ion runs. L.Betev agreed that ALICE should confirm it within a week.

Done.

  • 25 Nov 06 - Sites should send to H.Renshall their procurement plans.
Moved to End of January.

  • 19 Dec 2006 - K.Bos should distribute to the MB list the mandate, and participation, of the Storage Classes working group.
Done Proposal distributed. Feedback from the MB before next MB meeting.

  • 29 Nov 2006 - L.Robertson and F.Carminati will discuss with K.Bos about changes to the mandate of the Storage Classes working group.

Done. Discussed in the GDB meeting the previous week.

  • 21 Nov 06 - I.Bird will prepare the mandate, participation and goals of the working groups on Monitoring Tools and System Analysis.

Done. Mandates distributed. The message is in the MB mailing list archive (NICE password required).

  • 29 Nov 2006 – The MB member send feedback to A.Aimar on the Targets and Milestones for 2007.

Done. The Milestones will be discussed further in the ECM meeting but initial feedback was received.

  • 28 Nov 2007 - H.Renshall will send the updated table of the experiments requirements. These requirements are urgently needed by the sites in order to define their procurement plans.

Done. First version distributed, plans from the sites expected.

  • 13 Oct 2006 - Experiments should send to H.Renshall their resource requirements and work plans at all Tier-1 sites (cpu, disk, tape, network in and out, type of work) covering at least 2007Q1 and 2007Q2.

Done. H.Renshall presented the situation at the MB and will distribute the requirements.

  • 21 Nov 06 - F.Carminati agreed to prepare and propose (mail to the MB) the mandate, participation and goals of the working group on “Caching and Data Access”.
Done.

  • 27 Oct 2006 - The MB members should send to I.Bird names of candidates for coordination and participation to the three groups (Site Management, Monitoring and System Analysis).
Done, 14 Nove 06

  • 6 October - B.Panzer distributes to the MB a document on “where disk caches are needed in a Tier-1 site” everything included (buffers for tapes, network transfers, etc).

Done. B.Panzer distributed a document (document).

Comments and feedback are expected.

  • 10 Oct 2006 - The 3D Phase 2 sites should provide, in the next Quarterly Report, the 3D status and the time schedules for installations and tests of their 3D databases.

Done. The sites have included their in their plans the milestones for the installation of the 3D databases.

  • 31 Jul 06 - Experiments should express what they really need in terms of interoperability between EGEE and OSG. Experiments agreed to send information to J.Shiers.

Done.

D.Barberis distributed a mail with the ATLAS proposal for all grids interoperability (email).

  • 29 Sept 06 - MB members should send feedback and comments on the ECM, OPS and SCM meetings mandates and participation. Latest version

Done.

  • 6 Oct 2006 - J.Templon distributes information, as used at SARA, on how to calculate the disk cache size for the case disk0tape1.

Done.

  • 30 Jun 06 - J.Gordon reports on the defined use cases and policies for user-level accounting in agreement with the security policy working group, independently on the tools and technology used to implement it.

Done. J.Gordon presentd some use cases at the GDB in October

  • 12 Sept 2006 - H.Renshall and J.Shiers will distribute to the MB a summary with the mandate, input, output and participants of the OPS, SCM, RSM meetings.

Done.

  • 31 Aug 06 - Gonzalo Merino - to see if one of the Spanish Tier-2 federations will give a report on their status and difficulties at the Comprehensive Review.

Done.

  • 1 Sep 06 - Bernd Panzer - to organise a meeting with the experiment coordinators to review the effects of the revised estimates on the costs of the CERN facility.

Done.

  • 11 Sep 2006 - Les Robertson will propose to the Overview Board to show to the C-RRB the cpu, storage utilization, installed capacity, perhaps the pledge, with appropriate caveats, along with summary usage information by VO.

Done.

  • 15 Aug 06 - Sites should check the values of the resources availability and required and confirm them to H.Renshall. Provide explanations when there are fewer resources available than required.
To be removed. Will be replaced by a new table (and format) that will be distributed to the MB at next meeting.

  • 31 Jul 06 - Sites and experiments should check and update the data in the "Contact Page" page below.

From J.Shiers: The contacts page on the SC Wiki has been used for this purpose for >1 year and is regularly updated.

Done.

The VO information and contacts are always up to date on the CIC portal. I will modify the Wiki page to explain how to find the information (thanks to M.Barroso):

  1. Go to https://cic.in2p3.fr/index.php?id=vo
  2. Select the VO (alice, atlas, cms or lhcb) from the long list and click OK.
  3. All the information is shown (scroll down the long page)

  • 31 Jul 06 - Sites should exchange more information about monitoring, alarming and 24x7 support in the framework of HEPIX.

Done. Will be associates to the HEPIX meeting.

I.Bird proposed to HEPIX the organization of a workshop during next HEPIX meeting in October. For the moment it is not clear whether it will be possible. More news from I.Bird in the coming weeks.

J. Shiers notes that the SC Tech Day, to be held at CERN on September 15th 2006, has a session devoted to this issue.

  • 25 Aug 06 - All experiments - to provide by the end of the week (24 August) for each of their Tier-1 sites their requirements for T1-T1 bandwidth, T0-T1 bandwidth and storage space by mass storage class for purposes other than those already included in the T1/T2 relationship table (see email from Chris Eck).

Done. The revised numbers on experiments’ resource requirements for 2007 and 2008 were sent to C.Eck (CMS will do it before end of the week).

  • 15 Aug 06 - Experiments should provide a simple presentation of their monitoring information to the sites.
Cancelled.

  • 15 Aug 06 - L.Robertson will assemble the current estimates for energy, luminosity and efficiency for planning purposes, and distribute it to the MB list.
Done.

  • 23 May 06 - Tier-1 sites should confirm via email to J.Shiers that they have set-up and tested their FTS channels configuration for transfers from all Tier-1 and to/from Tier-2 sites. * Is it not sufficient to set up the channel but the action requires confirmation via email that transfers from all Tier-1 and to/from the "known" Tier-2 has been tested.

To be done: BNL, FNAL, NDGF and TRIUMF.

These sites will be asked individually to complete this action. The replies are in the MB minutes, here.

  • 15 Aug 06 - Experiments should confirm their participation to the MB in BNL (in person or by phone).

Done.

  • 20 Jul 06 - Tier-1 sites should nominate one or two representatives to the TCG. Nomination should be sent to I.Bird before the 18 July 2006

Done. The sites representatives at the TCG are Daniele Cesini and Jeff Templon in addition to Alessandra Forti. The MB approved the nominations.

  • 13 Jun 06 - D.Liko to distribute the Job Priority WG report to the MB.
Update. The report and a note on the status of the WG were distributed to the MB after the meeting.

Here it is:

http://egee-intranet.web.cern.ch/egee-intranet/NA1/TCG/wgs/priority.htm

Look for this section, and click

---------- 
First implementation plan (July 2006): 
    - Implementation plan 
    - WG Status 
----------

  • 31 May 06 - K.Bos should start a discussion forum to share experience and tools for monitoring rates and capacities and to provide information as needed by the VOs. The goal is then to make possible a central repository to store effective tape throughput monitoring information.

Action to be removed.

The fabrics and tools aspects could be followed up in the context of Hepix. I.Bird will ask the Hepix organizers. The service aspects are going to be discussed at the Service Challenge Technical Day, which is scheduled for the 15 September 2006.

  • 30 May 06 - ALICE will send to J.Shiers the list of the Tier-2 sites to monitor in SC4.
Done.

SC4 for LHCb does not include any Tier-2 sites. CMS list of Tier-2 sites: https://twiki.cern.ch/twiki/bin/view/CMS/SWIntSC4SiteStatus

  • 27 Jun 06 - K.Bos proposes a group that will look into implications of the move to SRM 2 for Tier-1 and Tier-2 sites.

Update from K.Bos' email to the GDB. The names of people that offered to take part in this group are:

  • Jan van Eldik (CERN)
  • Mark van de Sanden (SARA)
  • Artem Trunov (CMS and Alice, IN2P3)
  • Lionel Schwarz (IN2P3)
  • Adriŕ Casajús (PIC)
  • Jos van Wezel (FZK/GridKa)
  • Ruth Pordes as a place holder for possibly Eileen Berman and/or Frank Wuerthwein (FNAL)

  • 27 Jun 2006 - FZK sets up the ORACLE streams to CERN for the LCG 3D project.

  • 30 May 06 - J.Shiers and N.Brook will add a note in the document to remind of the need of announcing in advance draining of jobs at the sites.
Done. Final version to be distributed.

  • 16 Jun 06 - CERN+Tier-1 sites - The sites should send the accounting data for May to lcg.office@cern.ch

Done.

  • 20 Jun 06 - E.Laure prepares a clear statement on the status of support for MySQL in FTS
Done. Bug fixing will be done but no tests for high-performance usage of a mysql-based solution.

  • 10 Jun 06 - CMS will send to J.Shiers some defined milestones for the CMS SC4 exercises.

Done. The milestones are here: https://twiki.cern.ch/twiki/bin/view/CMS/SWIntSC4Mile

  • 10 Jun 06 � J.Shiers will add an Action List to the SC4 Plans wiki page.
Done. But there are not many clear actions.

  • 31 May 06 � C.Grandi presents to the MB the EGEE middleware priorities and development of the features needed by the LHC (in Flavia's list).

Done. This will be done in June.

  • 13 Jun 2006 - I.Bird to add the discussion on the SAM tests and results to the Operations Workshop agenda.

Done.

  • 15 June 06 � I.Bird reports to the TCG the needs of APIs to verify the working status of the middleware services.

Done.

  • 21 May 06 � J.Shiers and M.Schulz: Flavia�s list should be updated, maintained and used to control changes and releases. A fixed URL link should be provided to that list.

Done. The TCG list is the rference list and should be monitored carefully.

  • 28 Feb 06 - L.Robertson prepares a proposal to answer the recommendation on re-defining experiments support as an area will be distributed.

Cancelled.

  • 31 May 06 � J.Shiers proposes a plan for demonstrating capability to recover short and long interventions on the Tier-0 and Tier-1 sites.

Done. A paper for describing recovery from scheduled downtimes was distributed in May. The paper proposes that we first establish that we can run stable transfer services and then turn to the proposal to demonstrate recovery. The plan for testing such recovery will be produced later by the Service Coordination Meeting.

  • 31 May 06 � J.Shiers should clarify with the LCG Operations how, and to whom, the announcements of interventions should be distributed.
Done. See above.

  • 31 May 06 � J.Shiers define mechanism to coordinate maintenance interruptions of Tier-1s.
Done. And now there is also the verification that the procedures are followed by the sites.

  • 31 May 06 � M.Litmaath presents the results of the SRM workshop in FNAL.
Done at June's GDB.

  • 26 May 06 � Sites should complete and send the questionnaire for the Internal Review to the chair, V.Guelzow.
Update: Done. INFN questionnaire sent on the 7 June 2006.

  • 23 May 06 � Sites should send to lcg.office@cernNOSPAMPLEASE.ch the names of the contact persons attending, possibly in person, the Internal Review meeting.
Done. The Internal Review is on the 8-9 June 2006.

  • 20 May 2006 � All SC4 sites send accounting data using the report form that will be sent to them by Fabienne Baud-Lavigne (for Jan, Feb, Mar, Apr 2006).

Done. Accounting Data received. The summary is available and will be discussed at next MB meeting, feedback and suggestions are welcome.

The �Installed capacity� at CNAF is missing. It is needed in order to complete the statistics.

Update:

The accounting data for January to April 2006 is now available in:
http://lcg.web.cern.ch/LCG/MB/accounting/April2006_accounting_summaries.pdf
 
The Excel workbook used to generate the report is also available at the following web address:
http://lcg.web.cern.ch/LCG/MB/accounting/Master%20accounting%20summaries.xls
 
Please send me (or to lcg.office@cern.ch) any corrections and comments. 

  • 20 May 06 � Produce the agenda of the Internal Review of the LCG Services.
Update: Done. Here is the agenda: http://agenda.cern.ch/fullAgenda.php?ida=a062385

  • 20 May 06 � L.Robertson add a �comment� field to the site accounting sheet template.
Done.

  • 05 May 06 � L.Robertson will distribute a proposal on how to conclude the discussions on User Level Accounting.
Done.

  • 09 May 06 - L.Robertson will discuss with E.Laure and C.Grandi the status of the development of the features needed by the LHC (in Flavia�s list).
Done. L.Robertson discussed with C.Grandi about the features being developed in JRA1. C.Grandi will present priorities and next developments to the MB.

  • 23 May 06 � V.Guelzow will add a question on the �site monitoring system status� to the Internal Review�s questionnaire.
Action removed. It was decided by the reviewers that the questionnaire is already covering this issue.

  • 26 Apr 2006 B.Panzer - Detailed plan for the DAQ-T0-T1 activities in 2006 and match with the experiments milestones.

The outline of a first plan is described on this twiki page: https://cern.ch/twiki/bin/view/Main/DaqTierZeroTierOnePlanning
and is now under continuous update.

  • 1 Apr 06 - Sites should provide information to the GDB about any legal issue in their country concerming reporting of accounting data.
This is being done in the work on Accounting together with the security policy group.

  • Deployment group should define the sequence in which Tier-1s install gLite 3.0
Done. In the announce of the the gLite 3.0 release.

  • 15 Apr 06 � D.Duellmann should add performance targets and tests in the LCG 3D project plan.
Done. Added in the 2006Q1 the Quarterly Report. Now includes a proposed milestone for a dedicated throughput phase in May (hopefully this is inline with the request in the MB).

  • 25 Feb 06 � MB members and SC4 Contacts should send feedback and clarification to J.Shiers on the document about �Conclusions, Decisions and Open Issues after the Mumbai SC4 Workshop�.

Done. Feedback received and the updated document was sent to the GDB for comments and approval. * 28 Feb 06 - Experiments should clarify with B.Panzer the exact schedule (week no) and resources needed (KSI2000) in 2006.

On the way. A proposal with the available information will be presented next week during the MB face-to-face meeting.

  • 7 Mar 06 � J.Gordon reports to the MB about the feedback received on the accounting policies and to the request of installing accounting software.

  • 28 Feb 06 � The disk-tape tests should be re-instated in the SC3 throughput re-run at the beginning of the year, to be completed by the end of February. Such tests should verify that a steady satisfactory transfer rate can be reached and that past problems have been solved.

Done.

  • 28 Feb 2006 � J.Gordon will ask the sites to install the accounting software. If they don�t do it, they should explain the reasons for not doing so. J.Gordon will report to the MB early in March.

Done. J.Gordon mailed the MB, GDB and ROC managers about the accounting policy, asking for feedback. Summary and conclusions will be presented at the MB in March.

  • 31 Jan 06 - B.Panzer will discuss with the experiments and present to the MB a plan with possible dates and resources for the experiments and for IT activities (performance tests, etc). The plan will include a proposal for the ATLAS TDAQ large tests.
Done. A note has been written and presented to the MB members the meeting on the 7 Feb 2006.

  • 31 Jan 06 - L. Robertson will discuss with the areas and services managers in order to define measurable metrics for Phase 2.

Done. The metrics that will be adopted were presented by H.Renshall at the Mumbai SC4 workshop.

  • 31 Jan 06 - D.Foster should find information about the GEANT2 plans and send it to the MB. Needed to complete definition of milestone OPN-2

Done. Presented at the MB of the 7 Feb 2006. Update available at: http://lhcopn.cern.ch

  • 31 Jan 06 MB - J.Gordon prepares a presentation on the situation of grid accounting.

Done. Presented at MB (7 Feb 2006)

  • 20 Jan 06 - CNAF provide a plan for the deployment of CASTOR 2 at their site.
    • RAL: CASTOR 2 milestones are included in their general site milestones.
    • PIC: here is the CASTOR 2 plan.
    • ASGC: sent the link to their CASTOR 2 plan

Done. All plans are accessible from this page under the heading: Castor 2 Plans.

  • End 2005 - CMS provide a more �Tier-1 accessible� description of their models, data and workflow.

Done � draft provided on 11 February. The documents are all accessible from this page under the heading Simplified Computing Models.

  • End Jan 06 - L.Robertson: Organize phone meeting with the LHCC referees, end of January.

  • 15 Jan 06 - Areas, experiments and sites representatives complete the Quarterly Reports and send it to A.Aimar.

Done.

  • 17 Jan 06Tier-1 sites send updated information and dates about their SC3 tape tests. If they cannot do those tests in time for SC3 they must send the recovery plans in order to perform their tape tests as soon as possible and before the April SC4 throughput run.

Done. All sites have responded. Will participate: FZK, IN2P3, PIC, TRIUMF, SARA, BNL, DESY (Tier-2).

Most sites prefer to do the test re-run during the week starting the 30 January.

  • 13 Jan 06 � Experiments and sites should all send urgently the *name of one person with the authority for discussing the details of the plans for SC4).

Done.

  • 06 Dec 2005 - CNAF should send milestones for Tier1-Tier1 and Tier1-Tier2 operations. For next MB meeting.

Done.

See details at:

  • 06 Dec 2005 - ALICE should fill the VO boxes questionnaire on operations and send it to the GDB.
Done

  • 20 Dec 05 - D.Foster should find information about the GEANT2 plans and send it to the MB. Needed to complete definition of milestone OPN-2

Moved to end of January

  • 10 Jan 06 MB - J.Gordon prepares a presentation on the situation of grid accounting*, for the face to face MB meeting before the GDB (11 Jan 06).

The presentation was not scheduled because the MB meeting focused on other topics. Reschedule for the 31 January.

  • 6 Dec 05 - J.Shiers has initiated a team to define metrics to measure site reliability and availability, data transfer capability, and overall grid performance . He will distribute a proposal on this to the MB for input from the sites and services, before 6 December 2005.

On the way. Clarifications will take place during January. H.Renshall will produce the proposal and this issue is being discussed also in EGEE and several other groups. The metrics will clearly be refined and improved throughout the year based on experience.

  • 10 Dec 05 - Tier-1 Sites (at GDB) should complete their plans following the guidelines of the presentation given at the GDB. (GDB presentation).

On the way. Moved to middle of February because. Some sites said that they will send the updated plans when they have their 2006 plans discussed at the beginning of January.

  • 15 Jan 06 - A.Aimar will contact the people with pending actions and will update the action list.

Done.

  • 11 Jan 06 GDB - F. Rademakers: Presentation/demonstration of PROOF at the GDB in January, and preparation of the goals and plans for the evaluation.

There was a presentation about the Application Area in general by Pere Mato. This action is cancelled.

  • 20 Dec 05 - L.Robertson - A proposal to answer the recommendation on re-defining experiments support as an area will be distributed before end 2005.

Moved to End of February.

  • 10 Jan 06 - MB members propose candidates for the QuarterlyReports review team (one from a site one from an experiment).

Done. No candidates proposed.

  • 06 Dec 2005 - J.Shiers - Confirm the reporting and tracking of operations problems deriving from the CMS data transfer service.

Done. Several action have been done to avoid this problem in the future.
Move the LCG SC weekly conference call to Monday, right after the EGEE. Attendance should become more formal, with minutes of the meetings and with regular participation by sites and experiments. An action list should be maintained.

  • 06 Dec 2005 - J.Shiers and K.Bos will distribute a proposal on how experiments plan can be made available to Tier1 sites. For next MB.

Done.
A weekly report, based on existing reports in IT, will be distributed to the MB list.
The weekly reports currently available within IT are: (1) a weekly report of the EIS team on experiments activities, (2) the weekly report on LCG operations and (3) the EGEE operations report.

  • 12 Dec 05 - Sites should send updates to the capacity tables of the MoU.
A letter was sent to all centers asking for confirmation of: (1) the values in the MoU, (2) who will sign and (3) when they plan to sign the MoU. Sites should urgently reply.

Done.

  • 31 Dec 05 - A.Aimar will distribute the Quarterly Reports for 2005 Q4.
Done.

  • 10 Dec 05 - People responsible for each area and service should also define some measurable metrics to follow during Phase 2.
Replaced by an action (discuss with the sites and experiments) on L.Robertson for the 20 Jan 06.

  • 19 Dec 05 - L.Robertson should explain the Tier-1 milestones and the procurement issues at next Oversight Board meeting in December.

  • 20 Dec 05 - D.Kelsey - The current text in Section 1.5 should have reference mentioning that there will be another document explaining how the user data must be stored and how long must be kept.

  • 15 Dec 05 - The MB will review the situation of CASTOR on middle of December. When the plans for distributing Castor should have been defined.
A successful installation has been performed at CNAF. In January the same installation will be done in RAL. But all sites involved with Castor2 should provide their roll-out plan.

Action replaced by the action for the site to provide roll-out plans.

  • 6 Dec 05 - L.Robertson and A.Aimar will prepare the High Level Milestones for formal approval by the MB.

Done. And approved by the MB on the 20 Dec 2005.

  • 15 Dec 05 - Tier-1 Sites to set milestones on their known Tier-1-Tier-1 and Tier-1-Tier2 testing activities (IN2P3 and RAL stated specifically that they will do this).
    Update: RAL done. IN2P3 will do it in a few weeks

  • 06 Dec 2005 - FNAL should send milestones for Tier1-Tier1 and Tier1-Tier2 operations. For next MB meeting.

  • 06 Dec 2005 - CMS will send a contact name for the migration to Castor 2.

  • 6 Dec 05 - we need the final plan of the 3D project. D.Duellmann should contact the sites to explain what is being planned (including replication at Tier-1 and Tier-2 sites, and clarifying the usages of Oracle, Frontier, Squid, mysql, etc.)

  • 06 Dec 2005 - J.Shiers should circulate the full plan and tests for the SC3 throughput re-run. For next MB.

  • 06 Dec 2005 - J.Gordon will send more information to J.Shiers about the EGEE group on metrics.

  • 06 Dec 2005 - CNAF, NDGF, ASGC and FNAL should send feedback on the proposed targets for the SC3 re-run. For next week.

  • 29 Nov 05 - ATLAS will send a proposal about reporting on issues concerning non-EGEE resources to the MB.
  • 29 Nov 05 - CMS will send a similar statement to formally confirm the CMS opinion on OSG reporting.

Update: ATLAS and CMS have sent text and L.Robertson will integrate it and distribute it to the MB.

  • 29 Nov 05 - MB sends feedback on the GDB summary.
    Update: No feedback received. Security documents approved.

  • 29 Nov 05 - In order to make a list of action and react to the recommendations, the MB should look at the document and the LHCC reviewers� presentations. Comments expected before next meeting.

  • 22 Nov 05 - L.Robertson will check that the dates agreed (9 Dec and 10 Jan) for the workshop are suitable for C.Loumis.
    Update: The dates will be 24-25 January

  • L.Robertson will announce the decision to stop the Phase 2 Planning Group.

  • Tier-1 Sites to set milestones on their known Tier-1-Tier-1 and Tier-1-Tier2 testing activities (IN2P3 and RAL stated specifically that they will do this).
    RAL done.

  • All MB members to provide more feedback on the High Level Milestone is needed.
    Update: If no feedback comes next week, the High Level Milestones will be distributed for formal approval at next MB meeting.

-- AlbertoAimar - 09 May 2006

Edit | Attach | Watch | Print version | History: r6 < r5 < r4 < r3 < r2 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r6 - 2010-08-05 - DeniseHeagerty
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback